# An Information Theoretic Approach to Machine Unlearning

### v3

@@ 6.1 Benchmark Evaluation
Full-class Unlearning. We begin by comparing full class performance to that of the existing ZS methods. As seen in Tables 2 (a), 2 (b), and 4, JiT demonstrates significantly superior performance over GKT and EMMN, and is competitive with Boundary Shrinking. The authors of Chundawat et al. (2023b) note the poor scalability of both EMMN and GKT, which is evident in our results. Failing to scale to large problems or models is a significant barrier, since the value of unlearning is found mostly in large models that are expensive to train, or large datasets that are expensive to store. JiT performance is competitive with Boundary Shrinking despite having a fraction of the compute cost and, even compared to non-ZS SOTA, JiT performs reasonably; dropping only $0.6\%$ retain set performance compared to the baseline on ViT and outperforming both UNSIR and SSD on the MIA. Our performance also holds for VGG and, when using the same hyper-parameters for the face dataset, JiT generalizes well, outperforming SCRUB, SSD, and UNSIR.
@@ 8 Limitations
Since we minimise the gradient over each forget sample independently, we advise caution when using a model with batch normalization; since this changes the model’s mapping of single input/output to batch input/output (Gulrajani et al., 2017). This limitation can be mitigated by selecting an alternative choice of model or normalization strategy (e.g. layer norms).

### v4

@@ 6.1 Benchmark Evaluation
Full-class Unlearning. We begin by comparing full class performance to that of the existing ZS methods. As seen in Tables 2 (a), 2 (b), and 4, JiT demonstrates significantly superior performance over GKT and EMMN, and is competitive with Boundary Shrinking. The authors of chundawat2023zero note the poor scalability of both EMMN and GKT, which is evident in our results. Failing to scale to large problems or models is a significant barrier, since the value of unlearning is found mostly in large models that are expensive to train, or large datasets that are expensive to store. JiT performance is competitive with Boundary Shrinking despite having a fraction of the compute cost and, even compared to non-ZS SOTA, JiT performs reasonably; dropping only $0.6\%$ retain set performance compared to the baseline on ViT and outperforming both UNSIR and SSD on the MIA. Our performance also holds for VGG and, when using the same hyper-parameters for the face dataset, JiT generalizes well, outperforming SCRUB, SSD, and UNSIR.
@@ 8 Limitations
Since we minimise the gradient over each forget sample independently, we advise caution when using a model with batch normalization; since this changes the model’s mapping of single input/output to batch input/output (gulrajani2017improved). This limitation can be mitigated by selecting an alternative choice of model or normalization strategy (e.g. layer norms).